OpenStudy (anonymous):

integral of 1/(sqrt(36-100x^2))dx

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ du }{ \sqrt{a^2-u^2} } = \arcsin(\frac{ u }{ a }) +C\]

OpenStudy (anonymous):

does this help ?

OpenStudy (anonymous):

yes! im just confused with what to do with 100. Do i take it out of the integral?

OpenStudy (anonymous):

using this formula and let a = 6 u = 10x you see that a^2 = 36 u^2 = 100x^2 so it is exactly the form of the integral up there ..

OpenStudy (anonymous):

ok thank you!! I was making it more complicated then it needed to be

OpenStudy (anonymous):

yes .. :)